Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for March 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, March 31, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 24, 1999 to 7:00 a.m. Wednesday, March 31, 1999.

The Thermal Destruction Facility (TDF) was on-line for 159.5 hours and processed 4870.2 tons of contaminated soil during the time frame of this report. The TDF was on-line 94.9 % of the time with an average hourly feed rate of 30.5 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 276,310.4 tons of contaminated soil, which amounts to approximately 95.3% of the overall contaminated soil on-site.

During the time frame of this report, hazardous waste feed was stopped in order to perform maintenance to various TDF components. Feed was stopped briefly on Wednesday, March 24, 1999 in order to repair a faulty sensor that measures the rotational speed of the rotary kiln. Feed was stopped briefly on Thursday, March 25, 1999 in order to remove a piece of metal debris from the Dry Ash Drag (DAD) conveyor. Feed was stopped briefly on Tuesday, March 30, 1999 in order to repair a small gap that had developed between the floor plates of the DAD.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. ConCover was placed on the sides of the ash stockpile to help prevent erosion and fugitive emissions. Shot rock was imported from off-site to backfill the excavation floor for the 2-foot rock bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued south of the feed preparation building, adjacent to the CZ113 sheet piling. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. No drums or demolition debris were encountered. Once the materials were excavated, shot rock was backfilled against the sheet piling. The contractor continued to remove the soil on top of the HDPE liner inside of the feed preparation building. The soils were removed with a mini-excavator and hand held equipment in order to maintain the integrity of the liner.

During the time frame of this report, 2 roll-off containers of PPE, lime bags and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included:

* A weekly progress meeting was held on Tuesday, March 30, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* A site tour was given to students from Penn Technical College on March 26, 1999.

* A group of Army Officers from the Newcumberland Army Depot were given a tour of the site on March 30, 1999.

* The fourth and final stack testing event for metals and particulate was started on Tuesday, March 30, 1999.

* Kurt Davis, Technical Assistance Grant Representative, was on-site to observe stack testing on March 30, 1999.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, March 24, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 17, 1999 to 7:00 a.m. Wednesday, March 24, 1999.

The Thermal Destruction Facility (TDF) was on-line for 32.3 hours and processed 889.7 tons of contaminated soil during the time frame of this report. The TDF was on-line 19.2 % of the time with an average hourly feed rate of 27.5 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 271,831 tons of contaminated soil, which amounts to approximately 93.8% of the overall contaminated soil on-site.

Off-line time to the TDF during this reporting period is primarily attributed to the removal and repair of a damaged trunion bearing. The trunion bearing had a sheared shaft as reported in the previous weekly situation report. The bearing was rebuilt at a local machine shop and reinstalled. Hazardous waste feed to the incinerator restarted on Monday, March 22, 1999 at 1750 hours. Feed was later stopped at 1948 hours to repair a piece of ductwork, and the waste feed resumed at 0028 hours on Tuesday, March 23, 1999.

During the down time, OHM completed several maintenance tasks on the TDF. Maintenance activities included the removal of a buildup of material in the quench tower, the cleaning of the venturi nozzles in the quench tower, the cleaning of the fly ash conveyance system and the cleaning of the steam scrubber ductwork.

OHM's operations crews took advantage of the TDF down time to reorganize the equipment in the feed preparation building. The electric power grids and stacking conveyors that aid in the screening of debris were moved to allow excavation of the contaminated soils beneath the units. Once the soil was removed, the screeners and the conveyors were reassembled to process soil once the TDF was brought back on-line.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The utility corridor was regraded. Shot rock was imported from off-site to backfill the excavation floor for the 2-foot rock bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued in the vicinity of former sediment basin 2 directly south of the Waste Water Treatment Plant (WWTP) sheet piling. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. No drums or demolition debris were encountered. Once the materials were excavated, shot rock was backfilled against the sheet piling south of the WWTP.

During the time frame of this report, 4 roll-off containers of PPE and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included:

* A weekly progress meeting was held on Tuesday, March 23, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* A senior partnering meeting was held on Tuesday, March 23, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

* A site tour was given to students from Bucknell University on March 17, 1999.

* As a result of thermal operations being off-line last week, AirRecon did not perform stack testing during this reporting period. Stack testing is rescheduled for March 30 and 31, 1999.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, March 17, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 10, 1999 to 7:00 a.m. Wednesday, March 17, 1999.

The Thermal Destruction Facility (TDF) was on-line for 110.5 hours and processed 2905.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 65.8% of the time with an average hourly feed rate of 26.3 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 268,234 tons of contaminated soil, which amounts to approximately 92.6% of the overall contaminated soil on-site.

During the time frame of this report, hazardous waste feed to the incinerator stopped several times. Feed was manually stopped on Friday, March 12, 1999 in order to perform the monthly TDF calibrations and interlock checks (AWFSOs). During this down time, OHM performed preventative maintenance to both of the ash drags and cleaned residue from the inside of the steam scrubber duct work. Feed resumed during the evening hours on Saturday, March 13, 1999.

Feed to the incinerator was also manually stopped at 0927 hours on Tuesday, March 16, 1999 in order to replace a sprocket inside of the Wet Ash Drag (WAD). Feed was again stopped at 1653 hours when the kiln began rotating at a very low RPM rate. The kiln had slowed due to a problem with one of the trunion bearings, and a redundant drive motor was started to maintain RPMs. The contractor also retracted the dual feed screws in order to allow atmospheric air to cool the system. While performing the required lockout/tagout of the feed screws, electrical power to the Induced Draft (ID) fan was turned off, initiating a Thermal Relief Valve (TRV) opening at 1737 hours. Feed soil had been suspended for 44 minutes prior to the TRV opening which lasted for approximately 60 seconds. The ID fan was restarted, the TRV was closed, and the local authorities were notified of the event. Feed to the incinerator will not resume until the trunion bearing can be repaired and reinstalled.

The TDF experienced an additional (TRV) opening at 1240 hours on Thursday, March 11, 1999 when a new motor used in one of the TDF's air compressors caused a momentary power loss to the TDF. Power was quickly restored feed resumed at 1644 hours, once the unit reached operating temperatures. The proper local authorities were notified of the TRV opening.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. Shot rock was imported from off-site to backfill the excavation floor for the 2-foot rock bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued in the vicinity of sediment basin 2 directly south of the Waste Water Treatment Plant (WWTP) sheet piling. Much of the materials excavated consisted of fill soils and the naturally occurring soils at the bottom of the excavation. No drums or demolition debris were encountered. Excavation is also being performed on the inside of the feed preparation building as part of the site close out. The contaminated soil that was placed on top of the HDPE liner at the beginning of the project is being removed with a mini-excavator and hand equipment to protect the liner.

During the time frame of this report, 6 roll-off containers of PPE and miscellaneous site metal were shipped off-site as non-hazardous waste to a hazardous waste landfill. One roll-off containing decontaminated scrap metal was shipped off-site for recycling.

Related site activities included:

* A weekly progress meeting was held on Tuesday, March 16, 1999, with representatives from the City of Lock Haven, USEPA, PADEP, USACE, and OHM in attendance.

* On March 16, 1999, the USACE gave a site tour to representatives from the Local Emergency Planning Committee and a class of science students from Bucknell University.

* As a result of thermal operations being off-line, AirRecon will not be performing stack testing this week. Stack testing will be rescheduled for later this month.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, March 10, 1999

FOR DATES: From 7:00 a.m. Wednesday, March 3, 1999 to 7:00 a.m. Wednesday, March 10, 1999.

The Thermal Destruction Facility (TDF) was on-line for 148.8 hours and processed 3239.5 tons of contaminated soil during the time frame of this report. The TDF was on-line 88.6% of the time with an average hourly feed rate of 21.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 264,712.3 tons of contaminated soil, which amounts to approximately 91.4% of the overall contaminated soil on-site.

During the time frame of this report, hazardous waste feed to the incinerator was shut down for an extended period of time. Due to a local lightning storm, waste feed was shut down on March 4, 1999 at 1913 hours as part of the Standard Operating Procedure. A momentary power failure was experienced and, once the storm had passed, the burners were re-lit and the unit was brought back to operating temperature.

Due to the heavy rain and snow fall experienced during the early morning hours of Friday, March 5, 1999, power fluctuations caused two brief Thermal Relief Valve (TRV) openings. The TRV remained open for approximately 11/2 minutes each time. Hazardous waste feed to the incinerator had been off-line for approximately 12 hours prior to the openings, and the contractor notified the local authorities of both events. Additional off-line time was attributed to the removal of cemented ash from the inside of the double dump valve. Once the TRV was closed for the second time, the incinerator was brought up to operating temperature.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, wood management, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. Shot rock was imported from off-site to backfill the excavation floor for the 2-foot rock bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued in the vicinity of sediment basin 2 and directly south of the sediment basin. As part of the site close out plan, sediment basin 2 was eliminated and site storm water run-off has been diverted to large 22,000-gallon trailer containers called frac tanks. Much of the materials excavated recently consist of the soft saturated sediments that had accumulated in the sediment basin over the past year. The sediment was dewatered by mixing the soil with dry site soil and quicklime.

During the time frame of this report, two roll-off containers of PPE and miscellaneous metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included a weekly progress meeting held on Tuesday, March 9, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

On Friday, March 12, 1999, the TDF will be placed in 'hot hold' for approximately 8-hours in order to perform the monthly interlock checks (AWFSOs) and the monthly calibrations of the TDF equipment. During the down time, the thermal crews will be performing general preventative maintenance to the TDF equipment. As part of the interlock checks, a Thermal Relief Valve (TRV) opening will be initiated during the later morning hours. Prior to the TRV opening, feed to the incinerator will have been off for several hours and the proper local authorities will be notified.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan. The final quarterly stack-testing event for metals and particulate is scheduled to take place from March 17th to March 18th, 1999.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, March 3, 1999

FOR DATES: From 7:00 a.m. Wednesday, February 24, 1999 to 7:00 a.m. Wednesday, March 3, 1999.

The Thermal Destruction Facility (TDF) was on-line for 166.8 hours and processed 3981.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.3% of the time with an average hourly feed rate of 23.9 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 261,167 tons of contaminated soil, which amounts to approximately 90.4% of the overall contaminated soil on-site.

Site activities during the time frame of this report included TDF operations and maintenance, wastewater management, equipment decontamination, wood shredding, wood management, feed preparation, debris handling, debris separation, lime addition, ash handling, ash backfilling, dust control, excavation surveys, dry decontamination and backfilling of rock and concrete debris. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation activities were continued in the vicinity of sediment basin 2 and directly south of the sediment basin. As part of the project close out, the capacity of sediment basin 2 has been reduced. Site wastewater will now be stored in the reduced sediment basin and in large 22,000 gallon trailer containers. No drums were encountered during excavation activities.

During the time frame of this report, two roll-off containers of PPE and miscellaneous metal were shipped off-site as non-hazardous waste to a hazardous waste landfill.

Related site activities included a weekly progress meeting held on Tuesday, March 2, 1999, with representatives from the USEPA, PADEP, USACE, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling and drum bulking (if encountered), lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, rock backfilling, and the implementation of the project close out plan. The final quarterly stack-testing event is scheduled to take place from March 16th to March 18th, 1999.

PREPARED BY: Danny Zion for William Werntges (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.